OpenStudy (anonymous):

MEDAL!!!!!!!!! What is the solution to 5/x-3 +9=2? A. 16/7 B. 7/16 C. -7/16 D. -16/7

OpenStudy (anonymous):

@hartnn @xGuardians ?

OpenStudy (eric_d):

simplify it first 5/x-3= 2-9

OpenStudy (eric_d):

5=x-3(2-9)

geerky42 (geerky42):

\[\dfrac{5}{x-3}+9=2\]Correct?

OpenStudy (eric_d):

yup

OpenStudy (anonymous):

yes @geerky42

OpenStudy (eric_d):

5=x-3(-7) Expand it

OpenStudy (anonymous):

so -16/7?

OpenStudy (eric_d):

5=-7x+21 -16/-7=x x=16/7

OpenStudy (eric_d):

Got it

OpenStudy (anonymous):

Oh o thank you!!

OpenStudy (eric_d):

Wlcm